OpenStudy (anonymous):

An object is dropped from the top of a 100m tower. It's height above the ground after tsec is 100-4.9t^2m. How fast is it falling 2 seconds after it dropped?

OpenStudy (anonymous):

the m is just the meters notation, not meant to be part of the question as a variable

OpenStudy (baldymcgee6):

do you know calculus?

OpenStudy (anonymous):

take the derivative and replace \(t\) by 2

OpenStudy (anonymous):

i'm attempting to learn it, that's why I'm asking it

OpenStudy (anonymous):

so find the derivative, THEN replace t?

OpenStudy (anonymous):

oh crap you have to work from the definition ok not hard

OpenStudy (anonymous):

let me try that then get back to you @satellite73

OpenStudy (anonymous):

Again, take the derivative (haha....) and substitute 2. The derivative of a position function is speed.

OpenStudy (anonymous):

ok it is not hard or you can use \[\lim_{t\to 2}\frac{f(t)-f(2)}{t-2}\] if that would be easier

OpenStudy (anonymous):

I don't understand how you got that equation..

OpenStudy (anonymous):

it is another expression for \(f'(2)\)

OpenStudy (baldymcgee6):

Just take the derivative and dont worry about it then... if one way is easier, stick to it.

OpenStudy (anonymous):

She hasn't learnt the derivative formulas. Using the definition is VERY hard for this question.

OpenStudy (anonymous):

what @baldymcgee6 said. if you have one way, use it. but you are asked for the derivative at a number, so you can use \[\frac{f(2+h)-f(2)}{h}\] if you like, since you are working with \(2\) and not \(x\)

OpenStudy (anonymous):

It still looks very hard to simply all that with the definition...

OpenStudy (anonymous):

in the beginning it is probably easier to work with a number. you put a number in, get a number out

OpenStudy (anonymous):

is it 19.6 m/sec?

OpenStudy (anonymous):

it is the same as computing \[\frac{f(t+h)-f(t)}{h}\] and then replacing \(t\) by \(2\)

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

I see why you did the equation with replaced 2s, but it seemed to work thank God

OpenStudy (anonymous):

and I got it negative, but it's positive because it's speed? and that's always positive?

OpenStudy (anonymous):

So much easier with the power rule >.< .

OpenStudy (anonymous):

You should say the VELOCITY is positive.

OpenStudy (anonymous):

oops.. VELOCITY

OpenStudy (anonymous):

yes i think you are right, it is 18.6 in any case it is \(-2\times 4.9\times 2\)

OpenStudy (anonymous):

rest assured that in a week you will be able to do this annoying problem instantly in your head

OpenStudy (anonymous):

you will say " the derivative of \(100-4.9t^2\) is \(-9.8t\) and if you replace \(t\) by \(2\) you get \(-9.8\times 2\)"

OpenStudy (anonymous):

What satellite73 said. After you learn the derivative formulas it becomes MUCH easier.

OpenStudy (anonymous):

IT'LL BE THAT EASY? FML we're moving to a whole new chapter in the next week...

OpenStudy (anonymous):

It will be that easy :P . For instance I can tell you the derivative of e^tan(x) is sec^2 (x)*e^(tan(x) right away :P .